OpenStudy (anonymous):

sqrt property d^2-3=0

OpenStudy (anonymous):

sqrt3

jimthompson5910 (jim_thompson5910):

\[\large d^2-3=0\] \[\large d^2=3\] \[\large d=\pm\sqrt{3}\]

OpenStudy (anonymous):

thanks